OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6 A: slope= 5/3; y-intercept= (0, 2) B: slope= 5; y-intercept= (0, -6) C: slope= -3/5; y-intercept= (0, 2) D: slope= 5/3; y-intercept= (0, -6)

OpenStudy (anonymous):

5x-3y=-6 -3y=-5x-6 y=3/5x+2 slope=3/5 y-intercept=2

OpenStudy (anonymous):

Correction y=5/3x+2 slope= 5/3 y-int=2

OpenStudy (anonymous):

Thank you!(:

OpenStudy (anonymous):

you're welcome